OpenStudy (anonymous):

Suppose the temperature increases 8º to -7ºF. What was the starting temperature? a. 15º c. -13º b. -20º d. -15

OpenStudy (fibonaccichick666):

so you can write the eq x+8=-7 This defines this problem, can you solve it from here?

OpenStudy (anonymous):

- 15?

OpenStudy (fibonaccichick666):

That is correct. Do you know how I came up with the eq?

OpenStudy (anonymous):

x is the starting temperatute u have to add because it says it increases

OpenStudy (anonymous):

Add '-8' to each side of the equation. 8 + -8 + x = -7 + -8 Combine like terms: 8 + -8 = 0 0 + x = -7 + -8 x = -7 + -8 Combine like terms: -7 + -8 = -15 x = -15

OpenStudy (fibonaccichick666):

Yep, That is right. Any other questiions?

OpenStudy (anonymous):

yes 1 more

OpenStudy (anonymous):

When –3 is multiplied by the sum of a number and 8, the result is –21. Find the number. A. 3 B. –1 C. 8 D. –2

OpenStudy (fibonaccichick666):

Let's use the same approach as before, what eq can you make?